OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

(-2, 4) And (2, -2) y2 - y1 x2 - x1 4 - (-2) (-2) - 2 Solve.. 4 - (-2) = 6 (-2) - 2 = -4 Reduce 3 / -2